MARS DESERT RESEARCH STATION

Sol Summary – February 7th

Sol Summary:
Prepared by: Arpan Vasanth
Sol: 09
Earth date: 02.07.2017
The day began with regular team briefing session, we planned both today’s and tomorrow’s activities as we have a media person from Germany visiting us tomorrow! Team briefings is always fun, since its scheduled in the early mornings when the crew is full of energy.Today for the second time we had a break from cereals for the breakfast, our chief chef Sneha had prepared a typical Indian dish to kick start our day! In parallel Arun, Saroj and myself cleaned the vacuum cleaner to remove the martian dust!

I geared up with Saroj and Sneha for planned EVA after the breakfast, the goal was to get the soil samples from the farthest point from the hab. The team had finalised upon the location to be as Lith Canyon, it was very important to go to a region that had negligible or no human presence from previous missions as we need a ‘pristine’ sample of the desert soil. The EVA team buckled up for an exciting Dakar rally session except that the rovers was electrical! Its Mars folk’s renewable energy is very crucial. I was nevertheless very excited as Lith Canyon is via Burpee Dinosaur Quarry, I am fascinated by Dinosaurs since childhood and could not wait to see one of the Quarry in person! The site was magnificent with lots of information about different species of Dinosaur fossils found over there along with clear instructions of do’s and dont’s.

We looked around for a while and hit the road again driving through couple of tricky sessions wherein Sneha had to get off the vehicle to ensure safety! As we continued the route opened up to a mirage and we were mesmerised by the view, soon the camera was in action and we took a lot of pictures! Myself and Saroj went ahead with the protocols to sterilise the hands before collecting the samples, we found a lot of ‘alien’ footsteps so we had a walk a lot to finalise the spot! The mission was half done and with just 50 percent power left in one of the rover, we had to carefully evaluate and work our way back to the hab! Finally, after a long drive we could sense home after getting a glimpse of beautiful science dome! This was so far the longest EVA of mine and a memorable one.
Next activity was to check all the vital parameters for safe functioning of the hab for the operations report! There is nothing as satisfying as a warm meal after one is back from cold EVA, normally the crew in the hab takes responsibility for organising lunch. Today however it was a double treat as the hab pump was replaced by DG too! We have made it a point to have food together from the Sol0 and I feel it’s the best time there will be lots of interesting stories to share! Post lunch everyone started to work on the reports, while Sneha and Mamatha were occupied in the Greenhab watering the plants and monitoring root and shoot height of fenugreek plant! Later in the evening, our commander was busy sewing everyone’s mission patch onto suits as the Velcro was not holding up well! As mentioned earlier we are visited by media from Germany tomorrow and the team is looking forward for it!

GreenHab Report – February 7th

GreenHab Report
Sol: 09
Earth Date: 02.07.2017
Written by : Sneha Velayudhan

Functionality: The temperature in the green hab was noted to be 16 deg C with humidity of 30 %, 22.3 deg C with humidity of 35 % under the Misian Mars lamp while the outside temperature was 9 deg C at around 9:30 AM. The evening (5:30PM) temperature was noted as 21.8 deg C with 29% humidity, 24.7 deg C with 42% humidity under the Misian Mars Lamp while the outside temperature was 10 deg C. The heater is working fine.

Status: All plants were watered, the fenugreek seedlings were watered with vitamin D rich water and the growth of the seedlings were monitored and noted. The fenugreek seedlings under the Misian Mars lamp shows more growth than the seedlings in the green hab. The secondary roots are developing for most of the fenugreek seedlings growing under the Misian Mars lamp in Earth soil. There is an average of two leaves per plant and are dark green in color under the Misian Mars lamp. The ones in the green hab are growing slowly and steadily. The Mars sample soil pots in the green hab has dried out completely. Under the Misian Mars lamp, the Mars soil sample shows very less growth. The growth was noted and a graph was plotted as shown in figure below. The corn plants were measured and readings were noted.
Planned activity: The growth of the plants, temperature and other functionality will be monitored.

EVA Report – February 7th

EVA Report:
SOL: 09
Date: 02/07/2017
Report filed by: Saroj Kumar (Crew Scientist and Executive Officer)
Crew members involved in the EVA: Arpan Vasanth, Sneha Velayudhan and Saroj Kumar
EVA leader: Saroj Kumar
Begin: 10.45 am
End: 01.40 pm
Type of EVA: Walking + Rovers (Phobos and Deimos)
Purpose: Sample collection for characterising the transference of human commensal bacteria and developing zoning methodology for planetary protection and photography for outreach activities
UTM Coordinates: Northing: 4256715 Easting: 520914 Zone/Sector: 12S

Summary: The primary objective of today’s EVA was to collect sample number-7 for the experiment on developing zoning methodology for planetary protection. The first five samples for the experiment were taken at a distance of 0, 1, 2, 5 and 10 meters from the hab. Today’s samples were required to be collected from a very far distance from the hab with requirement of almost negligible or no prior human presence. Considering this, the crew selected ‘Green Mars north’ as one of the location for sample collection which is believed to have met the requirement. The secondary objective was towards outreach activities for which the crew had planned to take photographs of the sample collection process and take pictures of beautiful ‘Mars’ scenic view which looked similar to Valles Marineris.

At the beginning of EVA, the crew checked the battery level charge of rovers and made sure both Phobos and Deimos have sufficient charge to support the EVA. While heading towards ‘Green Mars north’ the crew spent few minutes at the entrance of ‘Special region’ to take some pictures. After reaching the destination, Arpan and Saroj took the samples while Sneha took the photographs of experiment procedure during sample collection. Later on, crew members took the pictures of beautiful Green Mars North. The crew spent about 45 minutes at the location and then headed back to the hab. During the end of the EVA after reaching hab, Arpan and Saroj took the readings and checked the status of hab support systems. The EVA for the day was successful by completing all the tasks as planned.